# GLOBE: Accurate and Generalizable PDE Surrogates using Domain-Inspired Architectures and Equivariances

### v2

@@ 1.4.3 Problem-Specific Constraints
Elliptic PDEs (e.g., steady-state aerodynamics, electrostatics, static structural analysis) exhibit global coupling: every point in the domain is influenced by every boundary condition and source term. Mathematically, well-posedness requires that boundary data at any location can affect the solution at any other location. For a surrogate model, this necessitates a global receptive field: the model’s output at any query point must have access to information from the entire domain.
Parabolic PDEs (e.g., heat diffusion, viscous flows) exhibit infinitely fast spatial information propagation, but distant influences decay rapidly with time. For small time steps $\Delta t$, local receptive fields can provide reasonable approximations; however, large-timestep solution operators still require spatially-distant information access.
@@ 4.1 Error Metrics
Model $\bar{u}_{x}\;(\times 10^{2})$ $\bar{u}_{y}\;(\times 10^{2})$ $\bar{p}\;(\times 10^{2})$ $\bar{p}_{s}\;(\times 10^{2})$
GLOBE (our work) 0.0047 0.0039 0.0031 0.0039
enf2enf 0.1100 0.3200
CORAL∗∗ 0.3500 1.2000
Transolver 0.3700 1.4200
MeshGraphNet†† 2.1400 3.8700
GINO|| 2.9700 4.8200
MLP 0.9490 0.9780 0.7370 11.3000
GraphSAGE 0.8320 0.9940 0.6610 6.6200
PointNet 3.5000 3.6450 1.1510 9.2500
GUNet 1.5170 2.0280 0.6570 3.8600
@@
Model $\bar{u}_{x}\;(\times 10^{2})$ $\bar{u}_{y}\;(\times 10^{2})$ $\bar{p}\;(\times 10^{2})$ $\bar{p}_{s}\;(\times 10^{2})$
GLOBE (our work) 0.0208 0.0180 0.0286 0.0392
MARIO§ 0.1520 0.1130 0.2400 2.7000
Transolver 2.1050 2.1080 4.4340 23.6700
MLP 1.6470 1.4510 3.9040 21.9200
GraphSAGE 1.4570 1.4540 4.6960 19.4500
PointNet 3.1110 2.7760 3.2940 18.2700
GUNet 1.7490 1.8250 3.3880 14.7300
@@
Model $\bar{u}_{x}\;(\times 10^{2})$ $\bar{u}_{y}\;(\times 10^{2})$ $\bar{p}\;(\times 10^{2})$ $\bar{p}_{s}\;(\times 10^{2})$
GLOBE (our work) 0.1897 0.0811 0.2446 0.2903
MLP 9.5050 4.9240 4.3000 208.9800
GraphSAGE 7.5580 3.4980 3.8260 17.9700
PointNet 9.4220 7.1290 4.0110 20.1300
GUNet 8.3830 5.2500 4.4830 20.5900
@@
Model $\bar{u}_{x}\;(\times 10^{2})$ $\bar{u}_{y}\;(\times 10^{2})$ $\bar{p}\;(\times 10^{2})$ $\bar{p}_{s}\;(\times 10^{2})$
GLOBE (our work) 0.9706 1.1928 11.3786 21.8215
MLP 6.9650 10.6300 11.7110 87.6200
GraphSAGE 4.4350 9.4000 10.9080 76.3800
PointNet 8.6800 15.7960 16.2370 58.4600
GUNet 5.6890 10.3420 14.8870 69.6700
@@ 4.3 Training and Inference on Non-Watertight Meshes
Model $\bar{u}_{x}\;(\times 10^{2})$ $\bar{u}_{y}\;(\times 10^{2})$ $\bar{p}\;(\times 10^{2})$ $\bar{p}_{s}\;(\times 10^{2})$
GLOBE (Full) 0.0047 0.0039 0.0031 0.0039
GLOBE (Full, decimated) 0.0510 0.0402 0.0311 0.0336
GLOBE (Scarce) 0.0208 0.0180 0.0286 0.0392

### v3

@@ 1.4.3 Problem-Specific Constraints
Elliptic PDEs (e.g., steady-state aerodynamics, electrostatics, static structural analysis) exhibit global coupling: every point in the domain is meaningfully influenced by every boundary condition and source term. Mathematically, well-posedness requires that boundary data at any location can affect the solution at any other location. For a surrogate model, this necessitates a global receptive field: the model’s output at a query point depends on information from the entire domain.
Parabolic PDEs (e.g., heat diffusion, most unsteady problems) exhibit infinitely fast spatial information propagation, but with rapid decay of influence with distance. For small time steps $\Delta t$, local receptive fields can provide reasonable approximations; however, large-timestep solution operators still require spatially-distant information access.
@@ 4.1 Error Metrics
Model $\bar{u}_{x}$ $\bar{u}_{y}$ $\bar{p}$ $\bar{p}_{s}$
GLOBE (our work) 0.0047 $\times 10^{-2}$ 0.0039 $\times 10^{-2}$ 0.0031 $\times 10^{-2}$ 0.0039 $\times 10^{-2}$
enf2enf 0.1100 $\times 10^{-2}$ 0.3200 $\times 10^{-2}$
CORAL∗∗ 0.3500 $\times 10^{-2}$ 1.2000 $\times 10^{-2}$
Transolver 0.3700 $\times 10^{-2}$ 1.4200 $\times 10^{-2}$
MeshGraphNet†† 2.1400 $\times 10^{-2}$ 3.8700 $\times 10^{-2}$
GINO|| 2.9700 $\times 10^{-2}$ 4.8200 $\times 10^{-2}$
MLP 0.9490 $\times 10^{-2}$ 0.9780 $\times 10^{-2}$ 0.7370 $\times 10^{-2}$ 11.3000 $\times 10^{-2}$
GraphSAGE 0.8320 $\times 10^{-2}$ 0.9940 $\times 10^{-2}$ 0.6610 $\times 10^{-2}$ 6.6200 $\times 10^{-2}$
PointNet 3.5000 $\times 10^{-2}$ 3.6450 $\times 10^{-2}$ 1.1510 $\times 10^{-2}$ 9.2500 $\times 10^{-2}$
GUNet 1.5170 $\times 10^{-2}$ 2.0280 $\times 10^{-2}$ 0.6570 $\times 10^{-2}$ 3.8600 $\times 10^{-2}$
@@
Model $\bar{u}_{x}$ $\bar{u}_{y}$ $\bar{p}$ $\bar{p}_{s}$
GLOBE (our work) 0.0208 $\times 10^{-2}$ 0.0180 $\times 10^{-2}$ 0.0286 $\times 10^{-2}$ 0.0392 $\times 10^{-2}$
MARIO§ 0.1520 $\times 10^{-2}$ 0.1130 $\times 10^{-2}$ 0.2400 $\times 10^{-2}$ 2.7000 $\times 10^{-2}$
Transolver 2.1050 $\times 10^{-2}$ 2.1080 $\times 10^{-2}$ 4.4340 $\times 10^{-2}$ 23.6700 $\times 10^{-2}$
MLP 1.6470 $\times 10^{-2}$ 1.4510 $\times 10^{-2}$ 3.9040 $\times 10^{-2}$ 21.9200 $\times 10^{-2}$
GraphSAGE 1.4570 $\times 10^{-2}$ 1.4540 $\times 10^{-2}$ 4.6960 $\times 10^{-2}$ 19.4500 $\times 10^{-2}$
PointNet 3.1110 $\times 10^{-2}$ 2.7760 $\times 10^{-2}$ 3.2940 $\times 10^{-2}$ 18.2700 $\times 10^{-2}$
GUNet 1.7490 $\times 10^{-2}$ 1.8250 $\times 10^{-2}$ 3.3880 $\times 10^{-2}$ 14.7300 $\times 10^{-2}$
@@
Model $\bar{u}_{x}$ $\bar{u}_{y}$ $\bar{p}$ $\bar{p}_{s}$
GLOBE (our work) 0.1897 $\times 10^{-2}$ 0.0811 $\times 10^{-2}$ 0.2446 $\times 10^{-2}$ 0.2903 $\times 10^{-2}$
MLP 9.5050 $\times 10^{-2}$ 4.9240 $\times 10^{-2}$ 4.3000 $\times 10^{-2}$ 208.9800 $\times 10^{-2}$
GraphSAGE 7.5580 $\times 10^{-2}$ 3.4980 $\times 10^{-2}$ 3.8260 $\times 10^{-2}$ 17.9700 $\times 10^{-2}$
PointNet 9.4220 $\times 10^{-2}$ 7.1290 $\times 10^{-2}$ 4.0110 $\times 10^{-2}$ 20.1300 $\times 10^{-2}$
GUNet 8.3830 $\times 10^{-2}$ 5.2500 $\times 10^{-2}$ 4.4830 $\times 10^{-2}$ 20.5900 $\times 10^{-2}$
@@
Model $\bar{u}_{x}$ $\bar{u}_{y}$ $\bar{p}$ $\bar{p}_{s}$
GLOBE (our work) 0.9706 $\times 10^{-2}$ 1.1928 $\times 10^{-2}$ 11.3786 $\times 10^{-2}$ 21.8215 $\times 10^{-2}$
MLP 6.9650 $\times 10^{-2}$ 10.6300 $\times 10^{-2}$ 11.7110 $\times 10^{-2}$ 87.6200 $\times 10^{-2}$
GraphSAGE 4.4350 $\times 10^{-2}$ 9.4000 $\times 10^{-2}$ 10.9080 $\times 10^{-2}$ 76.3800 $\times 10^{-2}$
PointNet 8.6800 $\times 10^{-2}$ 15.7960 $\times 10^{-2}$ 16.2370 $\times 10^{-2}$ 58.4600 $\times 10^{-2}$
GUNet 5.6890 $\times 10^{-2}$ 10.3420 $\times 10^{-2}$ 14.8870 $\times 10^{-2}$ 69.6700 $\times 10^{-2}$
@@ 4.3 Training and Inference on Non-Watertight Meshes
Model $\bar{u}_{x}$ $\bar{u}_{y}$ $\bar{p}$ $\bar{p}_{s}$
GLOBE (Full) 0.0047 $\times 10^{-2}$ 0.0039 $\times 10^{-2}$ 0.0031 $\times 10^{-2}$ 0.0039 $\times 10^{-2}$
GLOBE (Full, decimated) 0.0510 $\times 10^{-2}$ 0.0402 $\times 10^{-2}$ 0.0311 $\times 10^{-2}$ 0.0336 $\times 10^{-2}$
GLOBE (Scarce) 0.0208 $\times 10^{-2}$ 0.0180 $\times 10^{-2}$ 0.0286 $\times 10^{-2}$ 0.0392 $\times 10^{-2}$